OpenStudy (anonymous):

Juan has 3 coins. After Megan buys 5 more coins, how many more coins would Juan need to buy in order to have more coins than Megan? Let 'c' represent the number of coins that Juan needs to buy to have as many as Megan.

OpenStudy (anonymous):

How many coins did Megan have at the beginning? M(megan)+c=J(juan) 5+c=J

OpenStudy (anonymous):

\[3+c\text{ = }\text{megan}+5\]\[c=\text{megan}+2 \]